OpenStudy (anonymous):

PLEASE HELP will medal and fan The table below shows the ticket rates for whale watching trips offered by Ocean Tours: Age Ticket Price Under 3 years free 3 to 12 years $36 Over 12 years $48 On a certain day, the company took 92 people on whale watching trips. There were 40 children aged 12 and under, of which some children were under 3 years. If x represents the number of children under 3 years, which equation can be used to find the value of x, where C represents the total amount of money collected from tickets that day?

OpenStudy (anonymous):

A. (92 − x)36 + (92 − 40) ⋅ 48 = C B.(40 + x)36 + (92 − (40 + x)) ⋅ 48 = C C.(92 + x)36 + (92 − (40 + x)) ⋅ 48 = C D.(40 − x)36 + (92 − 40) ⋅ 48 = C

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

is it B?

jimthompson5910 (jim_thompson5910):

let me think

OpenStudy (anonymous):

well im pretty sure its b or a.. so

jimthompson5910 (jim_thompson5910):

x = number of children under 3 years so if "There were 40 children aged 12 and under", then you have 40-x children who are between 3 and 12, agreed?

OpenStudy (anonymous):

ohh so its D

OpenStudy (anonymous):

i read it wrong

jimthompson5910 (jim_thompson5910):

hold on, still thinking

jimthompson5910 (jim_thompson5910):

x = # of people under 3 40 - x = # of people age 3 to age 12 92 - 40 = 52 people ages 13 and up

jimthompson5910 (jim_thompson5910):

yeah so that makes it (40-x)*36 + (92 - 40)*48 in total dollars

OpenStudy (anonymous):

so D

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

yeah

jimthompson5910 (jim_thompson5910):

you're welcome